What is the molarity of a 250 mL solution containing 20 grams of KCl?


1
Expert's answer
2021-05-17T04:09:16-0400

v(solution)1=V(solution)2/Vm+m(KCl)/M(KCl)

Vm=22.4l/mol

v(solution)2=250ml/22400ml/mol=0.01116mol

M(KCl)=74.5g/mol

v(KCl)=20g/74.5g/mol=0.26846mol

v(solution)1=v(KCl)+v(solution)2=0.01116mol+0.26846mol=0,28mol